A car accelerates uniformly from rest to 20.0 m/s in 5.6 s along a level stretch of road. Ignoring friction, determine the average power required to accelerate the car if
(a) the weight of the car is 9.0 x 103 N

and

(b) the weight of the car is 1.4 x 104 N

Please show the reasoning and solution.
Answer #1
a) power=dW/dt=rate of change in ke with time= 1/2*m*v^2/t
=(1/2)*(9000/9.8)*20*20 / 5.6=35.714 kW
b)power=dW/dt=rate of change in ke with time= 1/2*m*v^2/t
=(1/2)*(1400/9.8)*20*20 / 5.6=5.102 kW